OpenStudy (owen3):

Find all solutions of (z+1)^5 = z^5. I am confused about something

OpenStudy (owen3):

I am looking at this solution here http://math.stackexchange.com/a/928749

OpenStudy (owen3):

How did they derive the following. 1. (z+1)^5 = z^5 implies | z + 1 | ^2 = |z|^2 2. | z + 1 | ^2 = |z|^2 implies Re(z) = - 1/2

OpenStudy (owen3):

1. take the fifth root ? then z +1 = z then |z + 1| = | z | 2. not sure

OpenStudy (owen3):

2. the distance from -1 is the same as the distance from 0. therefore the real part must be -1/2 .

OpenStudy (owen3):

| z - (-1) | = | z - 0 |

OpenStudy (owen3):